OpenStudy (anonymous):

y=-sin(x-pi/4)+2 amplitude, period, phase shift, vertical translation, graph the function I know the amplitude=-1 and vertical translation=up 2 but I don't know how to find period or phase shift

OpenStudy (anonymous):

Hi, \(\huge\color{yellow}{{Welcome~To~Open~Study}!!!!!!}\)

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

well amplitude is -1 vertical translation is up 2 right?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

or would amplitude be a positive number

OpenStudy (anonymous):

amplitude would be positive.. because it is distance so amplitude=1

OpenStudy (anonymous):

The standard form of a trig function is a*(trigfunction)(bx + c) + d

OpenStudy (phi):

see http://www.intmath.com/trigonometric-graphs/3-graphs-sin-cos-phase-shift.php for a quick summary

OpenStudy (anonymous):

a is amplitude for sin and cos b helps find the period (2pi/b = period) c is the phase shift d is the vertical shift

OpenStudy (anonymous):

so phase shift would be pi/4/1

OpenStudy (anonymous):

b is -pi/4 Period is (2pi)/(-pi/4)

OpenStudy (anonymous):

Whatever that simplifies to

OpenStudy (anonymous):

is it 8?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ y=-\sin\left(x-\frac{\pi}{4}\right)+2 \] match that up with \[ y = a \sin (bx + c) \] amplitude is the |a| or 1 (this means the sin moves up and down by 1 (total distance from the top to the bottom is 2) to find the period, match b with 2 pi/ T, where T is the period. b is 1 here, so \[ 1 = \frac{2 \pi}{T} \\ T = 2 \pi \] the period is 2 pi the phase shift is -c/b c is -pi/4 and b is 1 so you get phase shift = \( - \frac{- \pi}{4} = \frac{\pi}{4} \)

OpenStudy (anonymous):

ok. so for this problem: y=1/2cos(2x)-1 amplitude=1/2 vertical translation= down 1 period= pi phase shift=0 is this correct?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, your answers for the cos problem look good. btw, for the -1 * sin(x - pi/4) they may want us to rewrite it as 1 * sin (x - pi/4 + pi) which can be written sin(x + 3 pi/4) (in other words, the -1 can be thought of as causing a phase shift)

OpenStudy (anonymous):

thank you so much! and i need to graph these i can just put it in my calculator right? Y= -sin(x-pi/4)+2 and Y=1/2cos(2x)-1 and i enter it into the calculator just like that right?

OpenStudy (anonymous):

@phi